# First Comparative Exoplanetology Within a Transiting Multi-planet System: Comparing the atmospheres of V1298 Tau b and c

### v1

@@ 4.4 Comparing the early evolution of V1298 Tau b and c
We ran evolutionary models for the V1298 Tau planets. These models start from the time of disk dispersal, and model the structural and thermal evolution of the planet, including atmospheric mass loss due to photoevaporation (Owen & Wu, 2013). We run models assuming both a core-envelope interior structure, as well as with a gradually mixed interior. We have used the evolutionary models presented in Vazan et al. (2022). A brief description of the models is provided in Appendix G. Figure 2 shows a possible mass and radius evolution tracks for these planets.
@@
V1298 Tau c being closer to its host star compared to planet b, is more susceptible to atmospheric mass loss driven by photoevaporation. While the final mass and radius depend on photoevaporation efficiency and the star’s activity level (Poppenhaeger et al., 2021), our models show that V1298 Tau c is likely to lose a larger fraction of its current envelope compared to planet b. Therefore, even if both planets have similar masses now, their early evolution could lead them to different mass-radius regimes.
@@
Our models assume solar metallicity for the envelope opacity. However, the transmission spectrum of V1298 Tau c could also be explained using atmospheric hazes, unlike V1298 Tau b (Figure 1. Young planets are expected to cool and contract over time (Kubyshkina et al., 2020; Linder et al., 2019). The presence of high altitude hazes can increasing the atmospheric opacity, therefore increasing the cooling and contraction timescale (Lee & Chiang, 2015), leading to higher photoevaporative mass loss. Thus, a hazy atmosphere of V1298 Tau c as opposed to a clear atmosphere of V1298 Tau b, could lead to slower cooling and higher mass loss.
4.5 Comparing the composition of V1298 Tau planet with the exoplanet population
@@
Figure 3 shows V1298 Tau b and c on the mass-metallicity and metallicity-orbital distance planes. The observed low atmospheric metallicity for these planets deviates from the mass-metallicity trend predicted by core-accretion models (Pollack et al., 1996; Bodenheimer & Pollack, 1986) and reported for exoplanets (Welbanks et al., 2019; Thorngren & Fortney, 2019). This could potentially be explained by ongoing evolutionary processes that transform their current metal-poor atmospheres into the metal-rich atmospheres seen in mature sub-Neptune/super-Earth planets (Désert et al., 2011; Kempton et al., 2023; Libby-Roberts et al., 2020a; Madhusudhan et al., 2023).
@@
We compare the V1298 Tau system with the 30 Myr old HR 8799 system (Marois et al., 2010) with four directly imaged planets (HR8799 b, c , d and e) with masses 7, 10, 10, 10 (Marois et al., 2010; GRAVITY Collaboration et al., 2019). Nasedkin et al. (2024) using ground based spectra from the VLTI/GRAVITY program, report super-stellar(100 times stellar) atmospheric metallicity for these planets. Assuming a -0.65 dex metallicity for HR8799 (Swastik et al., 2021) we convert the retrieved atmospheric metallicity to stellar unit.
@@
The HR 8799 planets exhibit higher atmospheric metallicity compared to the mass-metallicity trend (Thorngren & Fortney, 2019) predicted by core-accretion formation, while V1298 Tau planets fall below this trend. HR 8799 is a system of widely separated massive gas giants, whereas V1298 Tau likely consists of compact, low-mass sub-Neptune/super-Earth progenitors. Both systems likely host primordial atmospheres. The different mass regimes suggest HR 8799 planets underwent runaway gas accretion or gravitational collapse, unlike V1298 Tau planets. Theories propose close-in sub-Neptunes and super-Earths formed in situ within the water ice-line in dusty or depleting disks to prevent runaway gas accretion (Lee et al., 2014; Lee & Chiang, 2016). Wide orbit gas giants likely formed beyond the ice lines, allowing for runaway accretion (Pollack et al., 1996) or gravitational instability (Boss, 1997; Dodson-Robinson et al., 2009), but these mechanisms alone do not explain HR 8799’s high metallicity, necessitating the accretion of volatile-rich solids (Nasedkin et al., 2024). Formation location relative to ice lines could lead to differences in volatile enrichment, as beyond the ice-line, the disk contains volatile-rich solids, enhancing the metal content of primordial atmospheres (Öberg et al., 2011; Bitsch et al., 2019).
@@ 4.6 Formation and evolution of multi-planet systems
The mass of the V1298 Tau planets indicates they are likely sub-Neptune/super-Earth progenitors, which are in a near mean-motion resonance (1:2:3). Atmospheric retrievals suggest these planets could have low metallicity atmospheres. Population synthesis studies (Mordasini, 2018; Benz et al., 2014) show sub-Neptunes and super-Earths often form in multi-planet systems in mean motion resonance (Lambrechts et al., 2019). Emsenhuber et al. (2023) classified systems with multiple low-mass planets into two types: Class I: inner dry/rocky planets formed in situ with outer icy sub-Neptunes migrating inwards, and Class II: multiple sub-Neptunes forming beyond the water ice line and migrating inwards. In the first scenario, inner planets have volatile-poor atmospheres, whereas, in the second, all planets are volatile-rich.
The relatively low atmospheric metallicity of V1298 Tau b and c suggests a formation similar to class I planets . However, these young planets’ atmospheric composition may evolve. In the pebble accretion scenario, a metallicity gradient forms in the protoplanet’s envelope (Ormel et al., 2021; Bodenheimer et al., 2018), with metal-rich layers hidden deep inside. Over time, atmospheric mass loss can strip volatile-poor upper layers, enriching atmospheric metallicity (Vazan et al., 2022), aligning them with mature sub-Neptune/super-Earth population. In this case, the V1298 Tau system could also have a formation similar to the class II systems.
If V1298 Tau c has a hazy atmosphere, it might be more metal-rich than the outer planet b, which doesn’t fit either class I or class II systems. Class I systems have inner volatile-poor planets, while class II systems have planets with similar compositions. Due to its proximity to the star, V1298 Tau c is likely to lose more of its primordial envelope compared to V1298 Tau b 2. Thus, V1298 Tau b and c might have formed like class II systems, with volatiles hidden deep inside, but atmospheric mass loss caused planet c’s composition to evolve faster than planet b, explaining their difference in their atmospheric composition.
@@ 5 Conclusion
In this paper, we perform the first comparative study of atmospheres of multiple planets in the same transiting system. Comparing the transmission spectra of the young V1298 Tau b and c planets reveals that planet b has a large, haze-free envelope, while the presence of hazes in planet c cannot be ruled out. Atmospheric models suggest similar masses for both planets in the sub-Neptune/super-Earth regime. These models also find relatively low atmospheric metallicities for both these young planets compared to their matured counterparts, indicating they are sub-Neptune/super-Earth progenitors. Alternatively, if planet c has a hazy atmosphere, higher haze formation in planet c compared to planet b could be due to differences in composition, temperature, and incoming stellar UV flux. The higher envelope opacity of planet c would affect its cooling rate and evolution differently than for planet b.
6 Acknowledgements

### v2

@@ 4.4 Comparing the early evolution of V1298 Tau b and c
We ran evolutionary models for V1298 Tau b and c. These models start from the time of disk dispersal, and model the structural and thermal evolution, including atmospheric mass loss due to photoevaporation (Owen & Wu 2013). We terminate atmospheric mass loss at 100 Myr. We run models assuming both a core-envelope interior structure, as well as with a gradually mixed interior. We have used the evolutionary models presented in Vazan et al. (2022). A brief description of the models is provided in Appendix D. Figure 4 shows a possible mass and radius evolution tracks for these planets.
We set up the initial conditions of our models such that these models can match the mass and radius of the planets at the current age (23 Myr). For planet c we assume 5.24R⊕ and 17 M⊕ (median value of mass from ATMO retrieval). For planet b we assume 9.95 R⊕ and show two models: a 10 M⊕ model which is consistent within 1 to the median value from our ATMO retrievals for planet b, and a 23 M⊕ model which is the 3 mass upper limit from Barat et al. (2024). We assume that the 10 M⊕ model should be similar to a 8 M⊕ model. For this calculation we do not account for the uncertainty in the mass of the planets, therefore, the models should be considered as potential evolutionary tracks.
Based on these models, V1298 Tau c could have started out with 20 M⊕ with a 25% by mass H/He envelope. Currently, it hosts 10% by mass H/He in its envelope and after 100 Myr of evolution it ends up as a Neptune mass planet with 1% by mass H/He envelope. After 100 Myr this planet enters into a phase of gradual thermal contraction and ends up with a radius between 2.5-3R⊕ (typical for Neptunes/sub-Neptunes). However, for V1298 Tau b, atmospheric mass loss is not as dominant, because it is further away from the host star, thereby, receiving 25% of the stellar flux compared to planet c. It starts off with 8.5 M⊕ and 23.5 M⊕ for the two models presented here, and loses less than 1 M⊕ in the course of its evolution. However, it is important to note here that Barat et al. (2024) report a high internal temperature (400K) for V1298 Tau b from disequilibrium chemistry models of its atmosphere, which has not been considered in our evolution models. Higher than expected internal temperature for planet b could lead to a puffed up atmosphere and enhanced atmospheric mass loss compared to our models.
The mass loss rates are strongly dependent on the current masses of the planets. In the scenario that planet b and/or c has mass lower than the estimates used for these calculations the H/He mass fraction in their envelope would be higher, leading to higher mass loss rates. In extreme scenarios, these planets could be completely stripped of their primordial atmospheres.
@@
While the final mass and radius depend on photoevaporation efficiency and the star’s activity level, from these models, we can conclude that due to its proximity to the star, V1298 Tau c is likely to lose a larger fraction of its H/He envelope compared to planet b. Therefore, even if both planets have similar masses now, their early evolution could lead them to different mass-radius regimes as well as on different ends of the radius valley.
@@
Our models assume solar metallicity for the envelope opacity. However, the transmission spectrum of V1298 Tau c could also be explained using atmospheric hazes, unlike V1298 Tau b (Figure 2. Young planets are expected to cool and contract over time (Kubyshkina et al. 2020; Linder et al. 2019). The presence of high altitude hazes can increase the atmospheric opacity, therefore increasing the cooling and contraction timescale (Lee & Chiang 2015), leading to higher photoevaporative mass loss. Thus, a hazy atmosphere of V1298 Tau c as opposed to a clear atmosphere of V1298 Tau b, could also lead to slower cooling and higher mass loss rates for planet c.
4.5 Comparing the atmospheres of V1298 Tau b and c with the exoplanet population
@@
Figure 5 shows V1298 Tau b and c on the mass-metallicity and metallicity-orbital distance planes. The observed low atmospheric metallicity (solar/sub-solar) for these planets, deviates from the mass-metallicity trend predicted by core-accretion models (Bodenheimer & Pollack 1986; Pollack et al. 1996) and reported for exoplanets (Welbanks et al. 2019; Thorngren & Fortney 2019). The atmospheric metallicities of planet b and c appears to be at least one order of magnitude lower compared to mature Neptune/sub-Neptunes. This could potentially be explained by ongoing evolutionary processes that transform their current metal-poor atmospheres of V1298 Tau b and c into the metal-rich atmospheres, as found in known mature Neptune/sub-Neptune planets (Désert et al. 2011a; Miller-Ricci Kempton et al. 2012; Morley et al. 2013; Libby-Roberts et al. 2020a; Bean et al. 2021; Kempton et al. 2023; Madhusudhan et al. 2023). One such possible mechanism has been suggested, where Neptune/sub-Neptunes are naturally born with metallicity gradients in their envelope (Ormel et al. 2021). Atmospheric mass loss during the early evolutionary phase removes the upper layers of the atmosphere, revealing the inner metal-rich layers (Fortney et al. 2013; Vazan et al. 2022). Thus, this mechanism could reconcile the the two Neptune/sub-Neptune progenitors, V1298 Tau b and c with their mature counterparts.
@@
We compare the V1298 Tau system with the 30 Myr old HR 8799 system (Marois et al. 2010) with four directly imaged planets (HR8799 b, c , d and e) with masses 7, 10, 10, 10 MJ, respectively (Marois et al. 2010; GRAVITY Collaboration et al. 2019). Nasedkin et al. (2024) using ground based spectra from the VLTI/GRAVITY program, report super-stellar(100 times stellar) atmospheric metallicity for these planets. Assuming a -0.65 dex metallicity for HR8799 (Swastik et al. 2021) we convert the retrieved atmospheric metallicity to stellar unit.
@@
The HR 8799 planets exhibit higher atmospheric metallicity compared to the mass-metallicity trend (Thorngren & Fortney 2019) predicted by core-accretion formation, while V1298 Tau planets fall below this trend. HR 8799 is a system of widely separated massive gas giants, whereas V1298 Tau likely consists of compact, low-mass Neptune/sub-Neptune progenitors. Both systems likely host primordial atmospheres. The different mass regimes suggest HR 8799 planets underwent runaway gas accretion (Pollack et al. 1996) or gravitational collapse (Boss 1997), unlike V1298 Tau planets. Formation theories of Neptune/sub-Neptunes suggest that these planets are born in situ within the water ice-line in dusty or depleting disks to prevent runaway gas accretion (Lee et al. 2014; Lee & Chiang 2016). Wide orbit gas giants likely formed beyond the ice lines, allowing for runaway accretion (Pollack et al. 1996) or gravitational instability (Boss 1997; Dodson-Robinson et al. 2009), but these mechanisms alone do not explain HR 8799’s high metallicity, necessitating the accretion of volatile-rich solids (Nasedkin et al. 2024). Formation location relative to ice lines could lead to differences in volatile enrichment, as beyond the ice-line, the disk contains volatile-rich solids, enhancing the metal content of primordial atmospheres (Öberg et al. 2011; Bitsch et al. 2019).
@@ 4.6 Formation and evolution of multi-planet systems
The mass of V1298 Tau b and c indicate that they are likely Neptune/sub-Neptune progenitors, which are in a near mean-motion resonance (1:2:3). Atmospheric retrievals suggest both these planets could have relatively low metallicity atmospheres. Population synthesis studies (Benz et al. 2014; Mordasini 2018) show that low mass planets, often form in multi-planet systems in mean motion resonance (Lambrechts et al. 2019). Emsenhuber et al. (2023) classified systems with multiple low-mass planets into two types: Class I: inner dry/rocky planets formed in situ with outer icy sub-Neptunes migrating inwards, and Class II: multiple sub-Neptunes forming beyond the water ice line and migrating inwards. In the first scenario, inner planets have metal-poor atmospheres, whereas, in the second, all planets are metal-rich.
The relatively low atmospheric metallicity of V1298 Tau b and c suggests a formation similar to class I planets . However, these young planets’ atmospheric metallicity may evolve. In the pebble accretion scenario, a metallicity gradient forms in the protoplanet’s envelope (Bodenheimer et al. 2018; Ormel et al. 2021), with metal-rich layers hidden deep inside. Over time, atmospheric mass loss can strip volatile-poor upper layers, revealing deeper metal-rich atmospheric layers(Fortney et al. 2013; Vazan et al. 2022), aligning them with mature sub-Neptune/super-Earth population. In this case, the V1298 Tau system could also have a formation similar to the class II systems.
If V1298 Tau c has a hazy atmosphere, it might be more metal-rich than the outer planet b, which doesn’t fit either class I or class II systems. Class I systems have inner volatile-poor planets, while class II systems have planets with similar compositions. Due to its proximity to the star, V1298 Tau c is likely to lose more of its primordial envelope compared to V1298 Tau b 4. Thus, V1298 Tau b and c might have formed like class II systems, with volatiles hidden deep inside, but atmospheric mass loss caused planet c’s metallicity to evolve faster than planet b.
@@ 5 Conclusion
In this paper, we perform the first comparative study of atmospheres of multiple planets in the same transiting multi-planet system. We compare the transmission spectra of two exoplanets in the young (20-30 Myr old) V1298 Tau system: V1298 Tau b and c. The transmission spectrum of V1298 Tau c shows an absorption feature that we interpret as due to water vapour (2.5), whereas, the spectra of its sibling, V1298 Tau b (adopted from Barat et al. (2024)), shows a clear atmosphere with a large scale height and a water absorption feature at 5 level of significance (Figure 2). We measure the masses of both planets from the transmission spectra ( M⊕ for planet b and M⊕ for planet c), together with the chemical abundances using a Bayesian chemical equilibrium retrieval framework using the ATMO atmospheric model. The retrieved mass for planet c is consistent with previous RV estimates (Suárez Mascareño et al. 2021; Sikora et al. 2023; Finociety et al. 2023). Our mass determination from the atmospheric scale heights also agree with ongoing TTV studies (Livingston et al, in prep). Our atmospheric models find sub-solar/solar atmospheric metallicity (logZ/Z⊙=-2.04 for planet b and logZ/Z⊙= -0.16 for planet c) for both these planets (Table 2), from free planet masses as well as fixed masses in the atmospheric retrievals. Compared to known mature Neptune/sub-Neptunes, which have been found to have metal-rich (100solar) and/or hazy atmospheres, the atmospheres of the two young Neptune/sub-Neptune progenitors (planet b and c), can be considered metal-poor by at least one order of magnitude. This hints at possible ongoing early evolutionary mechanisms which are expected to enhance the atmospheric metallicity of these planets, such that they are reconciled with their mature counterparts after evolution. Alternatively, the observed spectrum of V1298 Tau c could also be explained by atmospheric hazes. The large observed scale height of V1298 Tau b rules out high haze formation efficiency (, Figure 3) for this planet. Higher haze formation in planet c compared to planet b could be due to differences in bulk composition, temperature-pressure profile, and incoming stellar UV flux, which are known to be important factors affecting haze formation in H/He rich atmospheres (Kawashima & Ikoma 2018; Gao & Zhang 2020). In this scenario, the higher envelope opacity of planet c due to hazes could delay its cooling rate and compared to planet b, which has a clearer atmosphere. Evolutionary models suggest that planet c could lose a large fraction of its current H/He envelope (10% by mass) and end up with % by the end of it’s evolutionary phase. In contrast, our models applied to planet b, which receives 25% of the stellar flux received by planet c, do not predict significant mass loss and could end up retaining a substantial H/He envelope after 100 Myr of evolution.
Acknowledgements.